# Exponential separations between classical and quantum learners

### v1

@@ 1 Introduction
Quantum machine learning (QML) [${}^{+}$17, AdW17] is a bustling field with the potential to deliver quantum enhancements for practically relevant problems. An important goal of the community is to find practically relevant learning problems for which one can prove that quantum learners have an exponential advantage over classical learners. In this paper, we study how to achieve such exponential separations between classical and quantum learners for problems with classical data in the efficient probably approximately correct (PAC) learning framework. The results of this paper supersede the results of [GD22]. The first thing we address is that there is no single definition of what precisely constitutes a learning separation. In particular, when trying to come up with a definition there are many choices to be made, and various choices make sense depending on the particular settings. For instance, as we explain, a significant difference arise if the emphasis is on the task of identifying or evaluating the functions that are generating the data. This ambiguity can lead to conflating the task of learning in an intuitive sense with a purely computational task. To address this issue, we provide multiple definitions of a learning separation, and we discuss in which cases the tasks involve learning in an intuitive sense. Moreover, we study existing learning separations [LAT21, SG04] and carefully delineate where the classical hardness of learning lies and the types of learning separations they achieve. Furthermore, we provide new examples of learning separations where the classical hardness lies more in learning in an intuitive sense rather than evaluating the functions to be learned.
@@
  • We outline computational hardness assumptions that one can leverage to establish learning separations in the efficient PAC learning framework. In particular, we define the complexity class $\mathsf{HeurBPP/samp}$ that aims to capture all classically learnable functions (see Definition 13). Moreover, using ideas from [${}^{+}$21] we relate our new complexity class to a familiar though unexplored complexity class by showing that $\mathsf{HeurBPP/samp}\subseteq\mathsf{HeurP/poly}$ (see Lemma 2).
@@
  • We discuss known learning separations [LAT21, SG04], and we provide a fine-grained analysis of where the classical hardness of learning stems from. While discussing these learning separations, we find that the ones available in literature largely rely on the classical hardness of evaluating the function generating the data on unseen points, as opposed to the hardness of identifying it. We elaborate how the identification problem can be what is needed in practice, and we address this gap by proving two new learning separations where the classical hardness primarily lies in identifying the function generating the data (see Theorems 3 and 4).
@@
  • We show how leveraging stronger complexity-theoretic assumptions can lead to learning separations where the data is generated by a genuine quantum process. Our main contribution is Theorem 5, which outlines a method of establishing learning separations from $\mathsf{BQP}$-complete functions. We also provide two lemmas, Lemmas 3 and 4, which introduce natural assumptions under which the criteria in Theorem 5 are satisfied. Finally, we show how Theorem 5 can be used to build learning separations from problems in quantum many-body physics.
@@
  • To connect our work to some of the related results in the field [${}^{+}$22b, ${}^{+}$21, ${}^{+}$21, HKT21], we discuss selected topics related to learning separations with classical data:
@@
      We discuss the milestone work of Huang et al. [${}^{+}$22b] and how their classical machine learning methods based on the classical shadow framework relate to learning separations with quantum-generated data (i.e., those from Theorem 5). In particular, we highlight their limitations by constructing a family of Hamiltonians whose ground state properties cannot be predicted based on cryptographic assumptions (see Theorem 6).
@@ 2.1 Learning separations in the PAC learning framework
Definition 1 (Efficient PAC learnability).
A concept class $\mathcal{C}=\{\mathcal{C}_{n}\}_{n\in\mathbb{N}}$ is efficiently PAC learnable under target distributions $\mathcal{D}=\{\mathcal{D}_{n}\}_{n\in\mathbb{N}}$ if there exists a hypothesis class =$\mathcal{H}=\{\mathcal{H}_{n}\}_{n\in\mathbb{N}}$ and a (randomized) learning algorithm $\mathcal{A}$ with the following property: for every $c\in\mathcal{C}_{n}$, and for all $20<\epsilon<1/$ 2 and $20<\delta<1/$ 2, if $\mathcal{A}$ is given access to $EX(c,\mathcal{D}_{n})$ and $\epsilon$ and $\delta$, then with probability at least $1-\delta$, $\mathcal{A}$ outputs a specification [^2] The hypotheses (and concepts) are specified according to some enumeration $R:\cup_{n\in\mathbb{N}}\{0,1\}^{n}\rightarrow\cup_{n}\mathcal{H}_{n}$ (or, $\cup_{n}\mathcal{C}_{n}$) and by a “specification of $h\in\mathcal{H}_{n}$” we mean a string $\sigma\in\{0,1\}^{*}$ such that $R(\sigma)=h$ (see [KV94b] for more details). of some $h\in\mathcal{H}_{n}$ that satisfies
@@
In the above definition, the probability $1-\delta$ is over the random examples drawn from $EX(c,\mathcal{D}_{n})$ and over the internal randomization of $\mathcal{A}$. If the learning algorithm is a polynomial-time classical algorithm (or, a quantum algorithm), we say that the concept class is classically learnable (or, quantumly learnable, respectively). An important thing to note in the above definition is that the learner itself consists of two parts: a hypothesis class, and a learning algorithm. More precisely, the learner consists of a family of functions that it will use to approximate the concepts (i.e., the hypothesis class), and of a way to select which function from this family is the best approximation for a given concept (i.e., the learning algorithm). This is generally not very different from how supervised learning is done in practice. For example, in deep learning the hypothesis class consists of all functions realizable by a deep neural network with some given architecture, and the learning algorithm uses gradient descent to find the best hypothesis (i.e., the best assignment of weights).
@@ 2.2 Complexity theory
A distributional problem is a tuple $(L,\mathcal{D})$, where $L~{}\subseteq~{}\{0,1\}^{*}$ is a language [^4] Throughout this paper, we also use an equivalent definition of a language $L\subseteq\{0,1\}^{*}$ by instead calling it a problem and defining it as a function $L:\{0,1\}^{*}\rightarrow\{0,1\}$ such that $L(x)=1$ if and only if $x\in L$. and $\mathcal{D}=\{\mathcal{D}_{n}\}_{n\in\mathbb{N}}$ is a family of distributions such that $\mathrm{supp}(\mathcal{D}_{n})\subseteq\{0,1\}^{n}$.
@@
Note that in the PAC learning framework, the advice that the learning algorithm gets is of a specific form, namely that obtained through queries to the example oracle. This is more closely related to the notion of “sampling advice” complexity classes such as $\mathsf{BPP/samp}$ [${}^{+}$21] defined below. In [${}^{+}$21] it is shown that $\mathsf{BPP/samp}\subseteq\mathsf{P/poly}$, i.e., sampling advice is not more powerful than the standard notion of advice.
Definition 12 (Sampling advice [${}^{+}$21]).
@@
More precisely, if the concepts lie outside of $\mathsf{HeurBPP/samp}$, then the concept class is not classically learnable. We can connect the class $\mathsf{HeurBPP/samp}$ to other complexity classes by adopting a proof strategy similar to that of [${}^{+}$21] (we defer the proof to Appendix A.3).
@@
First proposed in Concepts based on Separation Complexity of concepts
@@
Table 1: The characteristics of the learning separations discussed in Sections 3 and 4.
@@ 3 Learning separations with efficient data generation
While discussing the learning separations of [LAT21, SG04], we notice that their proofs largely rely on the classical difficulty of evaluating the hypotheses on unseen examples, rather than the difficulty of identifying a hypothesis that is close to the concept generating the examples. To complement these works, we present two new examples of learning separations where the classical hardness lies in identifying the concept that is generating the examples. Specifically, in Section 3.3, we provide an example of a $\mathsf{CC/QC}$ separation (contingent on a plausible though relatively unexplored hardness assumption) where the concepts are classically efficiently evaluatable, making it impossible for the classical hardness to come from evaluating them on unseen examples. Afterwards, in Section 3.4, we provide an example of a separation in the setting where the learner is constrained to output hypotheses from a fixed hypothesis class, in which case the learner is only required to identify the concept generating the examples, therefore also eliminating the possibility that the classical hardness comes from evaluating them on unseen examples [^2] We remark that the concept class of Section 3.3 also exhibits a separation in the setting the learner is constrained to output a hypothesis from a fixed hypothesis class. However, we choose to present it as a $\mathsf{CC/QC}$ separation to highlight that such separations are still possible if the concepts are classically efficiently evaluatable. Moreover, we still include the separation in the setting the learner is constrained to output a hypothesis from a fixed hypothesis class of Section 3.4, because it is not contingent on a relatively unexplored hardness assumption..
3.1 A learning separation based on a worst-case to average-case reduction
In this section we discuss the discrete logarithm concept class studied in [LAT21]. In this work, the authors prove that the discrete logarithm concept class defined below exhibits a $\mathsf{CC/QQ}$ separation.
Definition 14 (Discrete logarithm concept class [LAT21]).
@@ 3.1 A learning separation based on a worst-case to average-case reduction
Fix an $n$-bit prime number $p$ and a generator $a$ of $\mathbb{Z}_{p}^{*}$ (i.e., the multiplicative group of integers modulo $p$). We define the discrete logarithm concept class as $\mathcal{C}^{\mathrm{DL}}_{n}=\{c_{i}\}_{i\in\mathbb{Z}^{*}_{p}}$, where
$c_{i}(x)=\begin{cases}1,&\text{if }\log_{a}x\in[i,i+\frac{p-3}{2}],\\ 0,&\text{otherwise.}\end{cases}$ (8)
Remark.
Here $\log_{a}x$ denotes the discrete logarithm of $x$ with respect to the generator $a$. That is, the discrete logarithm $\log_{a}x$ is the smallest positive integer $\ell$ such that $a^{\ell}\equiv x\mod p.$
@@
The weakest hardness assumption that one can leverage to obtain classical non-learnability is that the discrete logarithm is classically intractable (i.e., not in $\mathsf{BPP}$). Namely, in [BM84] it is shown that computing the most-significant bit of the discrete logarithm on any $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of inputs is at least as hard as computing the discrete logarithm on all inputs. Using the terminology of Section 2.2, if one assumes that the discrete logarithm is classically intractable (i.e., not in $\mathsf{BPP}$), then the concepts $c_{0}\in\mathcal{C}^{\mathrm{DL}}_{n}$ lie outside of $\mathsf{HeurBPP}$. In conclusion, to obtain the classical non-learnability it is sufficient to assume that the discrete logarithm is classically intractable.
We remark that one could obtain a similar learning separation for the singleton concept class $\mathcal{C}^{\prime}_{n}=\{c\}$ for any choice of $c\in\mathcal{C}^{\mathrm{DL}}_{n}$. This singleton concept class is quantum learnable without requiring use of the example oracle (i.e. without requiring any data), and one could thus argue that it is not a genuine learning problem anymore (i.e., similar to Observation 1).
Having discussed classical non-learnability, one still needs to ensure quantum learnability. To this end, the authors of [LAT21] show that a general-purpose quantum learning algorithm (i.e., a quantum kernel method) can efficiently learn the discrete logarithm concept class under the uniform distribution. We summarize the result of [LAT21] discussed in this section in the following theorem.
Theorem 1 ([LAT21]).
@@
$L_{\mathrm{DLP}}=\big{(}\{\mathcal{C}_{n}^{\mathrm{DLP}}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}})$ exhibits a $\mathsf{CC/QQ}$ separation, where $\mathcal{D}^{U}_{n}$ denotes the uniform distribution over $\mathbb{Z}_{p}^{*}$.
The hypothesis class the authors of [LAT21] use is quantumly evaluatable, and to the best of our knowledge it is unknown whether the discrete logarithm concepts are quantumly learnable using a classically evaluatable hypothesis class (which would imply a $\mathsf{CC/QC}$ separation).

### v2

@@ 1 Introduction
Quantum machine learning (QML) [BWP+17, AdW17] is a bustling field with the potential to deliver quantum enhancements for practically relevant problems. An important goal of the community is to find practically relevant learning problems for which one can prove that quantum learners have an exponential advantage over classical learners. In this paper, we study how to achieve such exponential separations between classical and quantum learners for problems with classical data in the efficient probably approximately correct (PAC) learning framework. The results of this paper supersede the results of [GD22]. The first thing we address is that there is no single definition of what precisely constitutes a learning separation. In particular, when trying to come up with a definition there are many choices to be made, and various choices make sense depending on the particular settings. For instance, as we explain, a significant difference arise if the emphasis is on the task of identifying or evaluating the functions that are generating the data. This ambiguity can lead to conflating the task of learning in an intuitive sense with a purely computational task. To address this issue, we provide multiple definitions of a learning separation, and we discuss in which cases the tasks involve learning in an intuitive sense. Moreover, we study existing learning separations [LAT21, SG04] and carefully delineate where the classical hardness of learning lies and the types of learning separations they achieve. Furthermore, we provide new examples of learning separations where the classical hardness lies more in learning in an intuitive sense rather than evaluating the functions to be learned.
@@
  • We outline computational hardness assumptions that one can leverage to establish learning separations in the efficient PAC learning framework. In particular, we define the complexity class $\mathsf{HeurBPP/samp}$ that aims to capture all classically learnable functions (see Definition 13). Moreover, using ideas from [HBM+21] we relate our new complexity class to a familiar though unexplored complexity class by showing that $\mathsf{HeurBPP/samp}\subseteq\mathsf{HeurP/poly}$ (see Lemma 2).
@@
  • We discuss known learning separations [LAT21, SG04], and we provide a fine-grained analysis of where the classical hardness of learning stems from.
    • We identify discrepancies between the definitions we posited in Section 2 and the definitions of the learning separations of [LAT21, SG04]. Upon noticing these discrepancies, we find that it is not directly clear how the learning separations of [LAT21, SG04] follow from the canonical hardness assumptions of the related computational problems (i.e., the discrete logarithm and discrete cube root outlined in [BM84] and [KV94b] respectively) when trying to make them comply with our definitions in Section 2. We address this as follows:
      • (i)
        We identify new stronger hardness assumptions for the relevant computational problems that adequately support a proof of learning separation according to our definitions in Section 2.
      • (ii)
        We introduce a new approach to translating the learning separations of [LAT21, SG04] to comply with our definitions in Section 2 by changing the underlying functions for which the hardness assumptions in [BM84] and [KV94b] are sufficient.
    • We find that the learning separations in literature largely rely on the classical hardness of evaluating the function generating the data, as opposed to the hardness of identifying the function. We discuss how the identification problem can be what is needed in practice, and we address this gap by proving two new learning separations where the classical hardness lies in identifying the function generating the data (see Theorems 7 and 8).
@@
  • We show how leveraging stronger complexity-theoretic assumptions can lead to learning separations where the data is generated by a genuine quantum process. Our main contribution is Theorem 9, which outlines a generic method of establishing learning separations from $\mathsf{BQP}$-complete functions. We also provide two lemmas, Lemmas 3 and 4, which introduce natural assumptions under which the criteria in Theorem 9 are satisfied. Finally, we show how Theorem 9 can be used to build learning separations from problems in quantum many-body physics.
@@
  • To connect our work to some of the related results in the field [HKT+22b, HBM+21, OIS+21, HKT21], we discuss selected topics related to learning separations with classical data:
@@
      We discuss the milestone work of Huang et al. [HKT+22b] and how their classical machine learning methods based on the classical shadow framework relate to learning separations with quantum-generated data (i.e., those from Theorem 9). In particular, we construct a family of Hamiltonians whose ground state properties cannot be predicted by any classical machine learning method based on cryptographic assumptions (see Theorem 10). These results show that the conditions needed to achieve learnability in [HKT+22b] essentially cannot be relaxed.
@@ 2.1 Learning separations in the PAC learning framework
Definition 1 (Efficient probably approximately correct learnability).
A concept class $\mathcal{C}=\{\mathcal{C}_{n}\}_{n\in\mathbb{N}}$ is efficiently PAC learnable under target distributions $\mathcal{D}=\{\mathcal{D}_{n}\}_{n\in\mathbb{N}}$ if there exists a hypothesis class =$\mathcal{H}=\{\mathcal{H}_{n}\}_{n\in\mathbb{N}}$ and a (randomized) learning algorithm $\mathcal{A}$ with the following property: for every $c\in\mathcal{C}_{n}$, and for all $20<\epsilon<1/$ 2 and $20<\delta<1/$ 2, if $\mathcal{A}$ is given access to $EX(c,\mathcal{D}_{n})$ and $(\epsilon,\delta)$, then with probability at least $1-\delta$ over the random examples drawn from $EX(c,\mathcal{D}_{n})$ and over the internal randomization of $\mathcal{A}$, the learning algorithm $\mathcal{A}$ outputs a specification [^2] The hypotheses (and concepts) are specified according to some enumeration $R:\cup_{n\in\mathbb{N}}\{0,1\}^{n}\rightarrow\cup_{n}\mathcal{H}_{n}$ (or, $\cup_{n}\mathcal{C}_{n}$) and by a “specification of $h\in\mathcal{H}_{n}$” we mean a string $\sigma\in\{0,1\}^{*}$ such that $R(\sigma)=h$ (see [KV94b] for more details). of some $h\in\mathcal{H}_{n}$ that satisfies
@@
Moreover, the learning algorithm $\mathcal{A}$ must run in time $\mathcal{O}(\mathrm{poly}(n,1/\epsilon,1/\delta))$. If the learning algorithm is a polynomial-time classical algorithm (or, a quantum algorithm), we say that the concept class is classically learnable (or, quantumly learnable, respectively).
An important thing to note in the above definition is that the learner itself consists of two parts: a hypothesis class, and a learning algorithm. More precisely, the learner consists of a family of functions that it will use to approximate the concepts (i.e., the hypothesis class), and of a way to select which function from this family is the best approximation for a given concept (i.e., the learning algorithm). This is generally not very different from how supervised learning is done in practice. For example, in deep learning the hypothesis class consists of all functions realizable by a deep neural network with some given architecture, and the learning algorithm uses gradient descent to find the best hypothesis (i.e., the best assignment of weights).
As we will discuss in more detail in Section 3, it is important to consider how the concept class is specified. What is important is that there must be some unambiguous definition of the concept class. In particular, the learnability criterion says that the concept class $\mathcal{C}$ is learnable if there exists an algorithm $\mathcal{A}_{\mathcal{C}}$ that satisfies the PAC criterion in Definition 1 with respect to $\mathcal{C}$. The point is that the algorithm $\mathcal{A}_{\mathcal{C}}$ is tailored to the specific concept class $\mathcal{C}$, and for this reason the concept class $\mathcal{C}$ has to somehow be unambiguously specified and fixed. For example, it is not acceptable that we leave some ambiguity as to what the concept class is precisely (e.g., it is specified by some sequence of primes, but you are not told which one), and then conclude that the concept class is not learnable because the learning algorithm would have to work for all possibilities that the ambiguity leaves open (which would be equivalent to a concept class involving all sequences of primes).
@@ 2.2 Complexity theory
A distributional problem $(L,\mathcal{D})$ consists of a language $L\subseteq\{0,1\}^{*}$ [^4] Throughout this paper, we also use an equivalent definition of a language $L\subseteq\{0,1\}^{*}$ by instead calling it a problem and defining it as a function $L:\{0,1\}^{*}\rightarrow\{0,1\}$ such that $L(x)=1$ if and only if $x\in L$. and a family of distributions $\mathcal{D}=\{\mathcal{D}_{n}\}_{n\in\mathbb{N}}$ such that $\mathrm{supp}(\mathcal{D}_{n})\subseteq\{0,1\}^{n}$.
@@
Note that in the PAC learning framework, the advice that the learning algorithm gets is of a specific form, namely that obtained through queries to the example oracle. This is more closely related to the notion of “sampling advice” complexity classes such as $\mathsf{BPP/samp}$ [HBM+21] defined below. In [HBM+21] it is shown that $\mathsf{BPP/samp}\subseteq\mathsf{P/poly}$, i.e., sampling advice is not more powerful than the standard notion of advice.
Definition 12 (Sampling advice [HBM+21]).
@@
More precisely, if the concepts lie outside of $\mathsf{HeurBPP/samp}$, then the concept class is not classically learnable. We can connect the class $\mathsf{HeurBPP/samp}$ to other complexity classes by adopting a proof strategy similar to that of [HBM+21] (we defer the proof to Appendix A.3).
@@
First proposed in Concepts based on Separation Complexity of concepts [^7] Assuming common assumptions in complexity theory and cryptography.
@@
Table 1: An overview of the characteristics of the learning separations discussed in Section 3 and Section 4. Importantly, in Section 4 we establish learning separations extending beyond the cryptographic problems discussed in Section 3 to encompass essentially all $\mathsf{BQP}$-complete problems.
@@ 3 Learning separations with efficient data generation
While discussing the learning separations of [LAT21, SG04], we will provide a fine grained analysis of the computational hardness assumptions necessary for establishing these separations. It is crucial to be precise about the specific computational hardness assumptions, as subtle details significantly impact the types of learning separations achievable. For instance, while the concept class based on the discrete logarithm from [LAT21] suggests a separation based on the canonical hardness assumption in [BM84], our fine-grained analysis reveals that this is not directly supported by the provided proof in [LAT21]. We address this by introducing a stronger modified hardness assumption and we show that this modification then suffices for a formal proof of classical non-learnability. Nevertheless, we will also show that similar learning separations are attainable assuming the standard hardness assumption in [BM84]. However, these assumptions require the introduction of a new set of concepts based on a novel construction, where information about the base of the discrete logarithm is “leaked” through examples in the dataset. Our detailed examination of the learning separation in [SG04] reveals a similar scenario, where we again could not recover a complete proof of classical non-learnability by leveraging the standard hardness assumptions in [KV94b], based on the details in [SG04]. We again address this issue in two ways: first, by introducing stronger hardness assumptions, which allow us to construct a proof of separation following the original approach of [SG04], and second, by introducing a new construction of the concept class, for which it is possible to prove hardness based on the standard hardness assumption in [KV94b].
In our fine-grained analysis, we also explore whether separations can be achieved for singleton concept classes, a crucial aspect in distinguishing genuine learning problems from problems that are just computational problems in disguise (as discussed in Section 2). We delve into how different hardness assumptions can lead to learning separations for singleton concept classes with different characteristics, revealing a trade-off. For example, while the computational hardness assumption we introduced to recover the learning separation of [LAT21] yields a $\mathsf{CC/QQ}$ separation for a binary singleton concept class, for the case of the typical hardness assumption in [BM84] we could only achieve a $\mathsf{CC/QQ}$ separation for multi-valued (i.e., non-binary) singleton concept class. Similarly, we note a trade-off in learning separations based on the discrete cube root assumption, as discussed in Sections 3.2-3.4. The computational hardness assumption introduced to recover the learning separation in [SG04] leads to a $\mathsf{CC/QC}$ separation for a singleton concept class, whereas our novel construction demonstrates that the typical hardness assumption in [BM84] results in a $\mathsf{CC/QQ}$ separation for a singleton concept class (i.e., a tradeoff in the type of separation). For an overview of these trade-offs, we refer to Table 3 & 3.
@@
Finally, while discussing the learning separations of [LAT21, SG04], we notice that their proofs largely rely on the classical difficulty of evaluating the hypotheses on unseen examples, rather than the difficulty of identifying a hypothesis that is close to the concept generating the examples. To complement these works, we present two new examples of learning separations where the classical hardness lies in identifying the concept that is generating the examples. Specifically, in Section 3.3, we provide an example of a $\mathsf{CC/QC}$ separation (contingent on a plausible though relatively unexplored hardness assumption) where the concepts are classically efficiently evaluatable, making it impossible for the classical hardness to come from evaluating them on unseen examples. Afterwards, in Section 3.4, we provide an example of a separation in the setting where the learner is constrained to output hypotheses from a fixed hypothesis class, in which case the learner is only required to identify the concept generating the examples, therefore also eliminating the possibility that the classical hardness comes from evaluating them on unseen examples [^2] We remark that the concept class of Section 3.3 also exhibits a separation in the setting the learner is constrained to output a hypothesis from a fixed hypothesis class. However, we choose to present it as a $\mathsf{CC/QC}$ separation to highlight that such separations are still possible if the concepts are classically efficiently evaluatable. Moreover, we still include the separation in the setting the learner is constrained to output a hypothesis from a fixed hypothesis class of Section 3.4, because it is not contingent on a relatively unexplored hardness assumption..
Concepts Properties Hardness assumption Separation Binary? Singleton?
Fix sequence $\{(p_{n},a_{n})\}_{n\in\mathbb{N}}$: $c_{i}(x)$ in Def. 14 $\mathsf{DLP}$-$\mathsf{fixed}$ $\mathsf{CC/QQ}$ Yes Yes
$c_{(a,p)}(x)$ in Def. 15 $\mathsf{DLP}$ $\mathsf{CC/QQ}$ Yes No
$c(x,a,p)$ in Def. 16 $\mathsf{DLP}$ $\mathsf{CC/QQ}$ No Yes
Table 2: An overview of the trade-offs regarding the types of separations achievable based on different hardness assumptions for concepts based on the discrete logarithm, providing a fine-grained analysis of the first row of Table 1. The hardness assumption $\mathsf{DLP}$-$\mathsf{fixed}$ states that there exists an efficiently generatable sequence for which the discrete logarithm is intractable, whereas $\mathsf{DLP}$ is the typical hardness assumption of the discrete logarithm outlined in [BM84]. We highlight that to achieve a learning separation for a singleton concept class based on the typical $\mathsf{DLP}$ assumption outlined in [BM84], the trade-off is that our construction has to consider non-binary concepts.
Concepts Properties Hardness assumption Separation Binary? Singleton?
Fix sequences $\{N_{i}\}_{i\in\mathbb{N}}$: $c_{j}(x)$ in Def. 17 $\mathsf{DCRA}$-$\mathsf{fixed}$ $\mathsf{CC/QC}$ Yes Yes
$c_{N,j}(x)$ in Def. 18 $\mathsf{DCRA}$ $\mathsf{CC/QC}$ Yes No
$c(N,x)$ in Def. 19 $\mathsf{DCRA}$ $\mathsf{CC/QQ}$ Yes Yes
Table 3: An overview of the trade-offs regarding the types of separations achievable based on different hardness assumptions for concepts based on the discrete cube root, providing a fine-grained analysis of the second row of Table 1. The hardness assumption $\mathsf{DCRA}$-$\mathsf{fixed}$ states that there exists an efficiently generatable sequence for which the discrete cube root is intractable, whereas $\mathsf{DCRA}$ is the typical hardness assumption of the discrete cube root outlined in [KV94b]. We highlight that to achieve a learning separation for a singleton concept class based on the typical $\mathsf{DCRA}$ assumption outlined in [BM84], the trade-off is that we achieve a $\mathsf{CC/QQ}$ instead of a $\mathsf{CC/QC}$ separation.
3.1 Learning separation based on heuristic hardness: case of discrete logarithm
In this section, we study learning separations for concept classes based on the discrete logarithm problem, as first studied in [LAT21]. The authors of [LAT21] introduce a concept class based on the discrete logarithm and they argue that it exhibits a learning separation which would correspond to a $\mathsf{CC/QQ}$ separation in our nomenclature. Additionally, they demonstrate its efficient learnability using a general-purpose quantum learning algorithm often referred to as a quantum kernel method.
Definition 14 (Discrete logarithm concept class).
@@ 3.1 Learning separation based on heuristic hardness: case of discrete logarithm
Let $\{p_{n}\}_{n\in\mathbb{N}}$ and $\{a_{n}\}_{n\in\mathbb{N}}$ be fixed sequences of $n$-bit prime numbers $p_{n}$ and generators $a_{n}$ of $\mathbb{Z}_{p_{n}}^{*}$ (i.e., the multiplicative group of integers modulo $p_{n}$). We define the discrete logarithm concept class as $\mathcal{C}^{\mathrm{DL}}_{n}=\{c_{i}\}_{i\in\mathbb{Z}^{*}_{p_{n}}}$, where
$c_{i}(x)=\begin{cases}1,&\text{if }\log_{(a_{n},p_{n})}x\in[i,i+\frac{p_{n}-3}{2}],\\ 0,&\text{otherwise.}\end{cases}$ (8)
Remark (Efficient data generation).
@@
Hardness assumptions
While [LAT21] suggests a separation for their concept class by leveraging the canonical hardness assumption of the discrete logarithm as outlined in [BM84], we note that we were unable to obtain a learning separation for the concept class in Definition 14 based on the hardness assumption in [BM84] by following the proof ideas in [LAT21]. Nonetheless, we will discuss two ways to address these shortcomings in the proof of separation. First, we propose introducing a new and stronger hardness assumption that can be used to obtain a full proof of a learning separation, while ensuring compatibility with quantum kernel methods (since we did not change the concepts when modifying the definition of the concept class). Next, we will explore an alternative approach to achieve a separation by changing the the concepts. This change will allows us to leverage the standard hardness of the discrete logarithm outlined in [BM84] to obtain a learning separation, albeit without direct compatibility with quantum kernel methods.
Our new hardness assumption (which we denote as $\mathsf{DLP}$-fixed) is that there exists a sequence of primes $\{p_{n}\}_{n\in\mathbb{N}}$ and generators $\{a_{n}\}_{n\in\mathbb{N}}$ for which the discrete logarithm is classically intractable. The modified result from [LAT21] is summarized in the following theorem.
Theorem 1 (modification of [LAT21]).
@@
$L_{\mathrm{DLP}}=\big{(}\{\mathcal{C}_{n}^{\mathrm{DLP}}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}})$ exhibits a $\mathsf{CC/QQ}$ separation assuming the intractability of the discrete logarithm problem for the fixed sequences of primes $\{p_{n}\}_{n\in\mathbb{N}}$ and generators $\{a_{n}\}_{n\in\mathbb{N}}$, where $\mathcal{D}^{U}_{n}$ denotes the uniform distribution over $\mathbb{Z}_{p}^{*}$.
Remark (Heuristic hardness).
A worst-case hardness assumption is sufficient for classical non-learnability, since the discrete logarithm admits a worst-to-average case reduction [BM84]. Note that this worst-to-average case reduction holds for both the canonical hardness assumption as outlined in [BM84] as well as our newly introduced and stronger $\mathsf{DLP}$-fixed assumption.
However, we note that the $\mathsf{DLP}$-fixed assumption is stronger than the typical one concerning the discrete logarithm, as outlined in [BM84]. In particular, the hardness assumption outlined in [BM84] states that there does not exist a polynomial-time classical algorithm that can compute the discrete logarithm for most primes and generators, not that there does not exist a polynomial time algorithm for any fixed sequence of primes and generators. It would thus be preferable to work with the weakest assumption and construct learning separations that leverage the canonical hardness of the discrete logarithm outlined in [BM84]. We show that this is indeed possible and that one can relax the requirement for the existence of sequences of primes and generators for which the discrete logarithm is classically intractable. Specifically, one can define a different yet closely related concept class, that uses a different input space and underlying set of functions, and show that it exhibits a $\mathsf{CC/QQ}$ separation assuming the canonical hardness of the discrete logarithm outlined in [BM84].
Definition 15.
We define the concept class
$\hat{\mathcal{C}}^{\mathrm{DL}}_{n}=\{c_{(a,p)}\mid p\text{ an $n$-bit prime and }a\in\{1,\dots,p-1\}\},$
where
$c_{(a,p)}(b,j,x)=\begin{cases}\mathds{1}_{[0,(p-1)/2]}(\log_{(a,p)}(x))\quad\text{if }b=0\\ \mathrm{bin}((a,p),j)\quad\text{if }b=1.\end{cases}$
for $b\in\{0,1\}$, $j\in[2n]=\{1,\dots,2n\}$ and $x\in\{0,1\}^{n}$.
This new concept class essentially involves a mechanism for leaking information about the prime and generator forming the base of the logarithm of the concept in the data. We delve deeper into this construction in Appendix B.1 (which may be of independent interest in other contexts) where we provide a formal proof of the following theorem.
Theorem 2.
$\hat{L}_{\mathrm{DLP}}=\left(\{\hat{\mathcal{C}}^{\mathrm{DL}}_{n}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}}\right)$ exhibits a $\mathsf{CC/QQ}$ separation assuming the intractability of the discrete logarithm problem as outlined in [BM84], where $\mathcal{D}_{n}^{U}$ denotes the uniform distribution over $\mathcal{X}_{n}=\{0,1\}\times[2n]\times\{0,1\}^{n}$.
Singleton concept class separations
Another question we would like to address is whether separations are possible for singleton concept classes. As discussed in Section 2, this is crucial in determining whether a problem is a genuine learning problem or merely a computational problem in disguise. Firstly, we observe that a learning separation for the singleton concept class $\mathcal{C}^{\prime}_{n}=\{c\}$, for any choice of $c\in\mathcal{C}^{\mathrm{DL}}_{n}$ from Definition 14, can be achieved if one assumes our stronger $\mathsf{DLP}$-fixed hardness assumption. However, one might wonder whether it is also possible to get a separation for a singleton concept class assuming the canonical hardness assumption of [BM84]. It turns out that this is indeed possible, though it again requires one to define a different yet closely related singleton concept class that utilizes a similar construction as those in Definition 15.
Definition 16.
We define the concept class $\hat{\mathcal{C}}^{\mathrm{DL}}_{n}=\{c_{n}\}$, where
$c_{n}(a,p,x)=\log_{(a,p)}(x),\quad\text{for $a,p,x\in\{0,1\}^{n}$.}$
Theorem 3.
$\hat{L}_{\mathrm{DLP}}=\left(\{\hat{\mathcal{C}}^{\mathrm{DL}}_{n}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}}\right)$ exhibits a $\mathsf{CC/QQ}$ separation assuming the intractability of the discrete logarithm problem as outlined in [BM84], where $\mathcal{D}_{n}^{U}$ denotes the uniform distribution over $\mathcal{X}_{n}=\{(a,p,x)\mid p\text{ an $n$-bit prime},a\in\{0,\dots,p-1\}\text{ and }x\in\mathbb{Z}^{*}_{p}\}$.
A full proof of the above theorem can be found in Appendix B.2. We note that in our construction the learning separation for a singleton concept class assuming the the canonical hardness assumption of [BM84] comes at the expense of making the concepts have non-binary labels. We thus notice a trade-off where stronger assumptions are required for separations for binary singleton concept classes, while the difficulty assumption can be relaxed if we consider non-binary concepts.